What average mechanical power (in W) must a 61.5 kg mountain climber generate to climb to the summit of a hill of height 345 m in 49.0 min? Note: Due to inefficiencies in converting chemical energy to mechanical energy, the amount calculated here is only a fraction of the power that must be produced by the climber's body.


__W


1
Expert's answer
2021-09-26T18:48:53-0400

The power is work over time:


"P=\frac Wt=\frac{mgh}t=\frac{61.5·9.8·345}{49·60}=70.7\text{ W}."